OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of . Explain in your own words how you can simplify without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification. x^4 ----- Expanded Form. x^2 This is equal to x^2, as when dividing powers you subtract the power of the denumerator from that of the numerator. 4 - 2 = 2, therefore the power of the resultant is 2. This is how you can know the power of the simplified form without writing the expanded. Part 2: x^7 ---- = x^4 x^3 Does this make sense,

OpenStudy (anonymous):

Is it correct?

Directrix (directrix):

I'm trying to understand what the problem is. That is, what is the question? I don't know what is being expanded.

OpenStudy (anonymous):

This is the question: Part 1: Show the expanded form and simplification of . Explain in your own words how you can simplify without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

x^4/x^2

OpenStudy (anonymous):

Sorry about that...

Directrix (directrix):

No reason to be sorry. Show the expanded form and simplification of --> ** what comes after the "of" **? Is it x^4/x^2 or something else.?

OpenStudy (anonymous):

yeah, its x^4/x^2

Directrix (directrix):

Part One x^4/x^2 = [ x * x * x * x ] / [ x * x] = x/x * x/x * x * x = 1 * 1 * x * x = x^2

Directrix (directrix):

Is "coefficients" supposed to be "exponents?"

OpenStudy (anonymous):

Ohh wow, thanks for breaking it down like that! :) Does my wording sound good though?

OpenStudy (anonymous):

In the question?

Directrix (directrix):

Yes

OpenStudy (anonymous):

Its given as coeficients

OpenStudy (anonymous):

So no its not suppose to be exponents.

OpenStudy (anonymous):

So, does the wording I have sound good, or should I change it?

OpenStudy (anonymous):

and is part 2 okay?

Directrix (directrix):

This is equal to x^2. When dividing exponential expressions with the same base, you subtract the power of the denumerator from that of the numerator. 4 - 2 = 2. The result is 2. This is how you can know the power of the simplified form without writing the expanded.

OpenStudy (anonymous):

Ohh thanks so much!! :)

Directrix (directrix):

Where did you do part 2? What is your fraction? Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

I have either x^7 ---- = x^4 x^3 or x^9 ---- = x^5 x^4

Directrix (directrix):

You don't have any coefficients. :)

Directrix (directrix):

4 x^9 ---- = 2 x^5 2 x^4

Directrix (directrix):

4 and 2 are coefficients.

OpenStudy (anonymous):

Ohhh wow, I can't believe I missed that one... lol :) Thanks so much for all of your help! I suck at math, so Its cool to have someone to break it down and explain it. :D

Directrix (directrix):

Okay. It seems to be that you are doing good. Don't say you suck at math. You might believe it. :)

OpenStudy (anonymous):

Yeah, thanks again! lol :)